Suppose (e1,e2,...,en) is an orthonormal basis of the inner product space V and v1,v2,...,vn are vectors of V such that ||ej−vj||<1/√n. Prove that (v1,v2,...,vn) is a basis of V.


1
Expert's answer
2021-07-28T17:37:34-0400

"\text{Note: It is enough to show that \((v_1 \ldots, v_n)\) are linearly independent.}\\
\text{Suppose \(v_i\) are linearly dependent then there are scalars \((\lambda_1 ,\ldots ,\lambda_n)\) whic are not all zero such that } \\
\sum_{i=1}^{n} \lambda_i v_i = 0 \\
\text{Then }\\
 0 = \sum_{i=1}^{n} \lambda_i v_i = \sum_{i=1}^{n} \lambda_i (v_i - e_i) + \sum_{i=1}^{n} \lambda_i e_i \\
\implies \| \sum_{i=1}^{n} \lambda_i (v_i - e_i ) \|= \| \sum_{i=1}^{n} \lambda_i e_i \| \\
\text{However, } \\
\|\sum_{i=1}^{n} \lambda_i e_i\| =\|\sum_{i=1}^{n} \lambda_i( v_i-e_i) \| \le \sum_{i=1}^{n} |\lambda_i| \|v_i - e_i\| \\
\qquad\qquad\quad\quad < \frac{1}{\sqrt{n}} \sum_{i=1}^{n} |\lambda_i| \le \left(\sum_{i=1}^{n} |\lambda_i|^2 \right)^{\frac{1}{2}} \,\,\,\,\text{(Cauchy-Schwarz)} \\

\implies \|\sum_{i=1}^{n} \lambda_i e_i\| \le \left(\sum_{i=1}^{n} |\lambda_i|^2 \right)^{\frac{1}{2}} \,\,\,\, (1)\\
\text{Since \(e_i\) are orthonormal we have }\\
\|\sum_{i=1}^{n} \lambda_i e_i\|^2 = \sum_{i=1}^{n} |\lambda_i|^2 \\
\text{So we have: }\\
\|\sum_{i=1}^{n} \lambda_i e_i\| = \left(\sum_{i=1}^{n} |\lambda_i|^2 \right)^{\frac{1}{2}}\\
\text{which is a contradiction to what we have above} \\\text{ (i.e inequality (1)) hence showing that \((v_1, \ldots,v_n)\) are linearly indepemdent and we have our result.}"